This page contains two parts. In the Timeline of instrument status all events affecting data quality and calibrations are listed. Announcements list current and past news items on instrument modes and operation.

Timeline of instrument status

The following table describes the events affecting the quality or the measurement of data obtained with UVES. Its content can be downloaded as a Comma Separated Values file. It is complete as of November 21, 2021.

The Start Date column indicates when the event took place. Some events cannot be defined by a single instant, but extend over sometimes several days, weeks or months. Operation with the instrument may continue while the event occurs, in which case the Ongoing flag is set. If the event had ended, the Ongoing ? flag is lowered and End Date is given. The Type column indicates the characteristic(s) of the event: Earthquake, Maintenance, Upgrade, Element failure, Intervention, Pipeline installation at QC, Change of nominal values or thresholds, Modification of configuration file of this database, or Other. The Instrument Modes column indicates the affected mode. The column KPI indicates which quality control parameters identified as Key Performance Indicators (i.e., parameters following the most critical quantities) are affected; the column non-KPI indicates which other critical Quality Control parameters are affected. The Calibration Raw Types column indicates the raw types of the affected calibrations. The Calibration Breakpoint flag indicates if calibrations obtained before the event cannot be used for science data obtained after the event or if calibration data obtained after the event cannot be used with science data obtained before the event. The Comment column provides more details. The History column allows one to see possible past information on the entry.

The Quality Control pages for UVES can be accessed through the QC overview page.



  • The procedure for taking flux standard for the purpose of determining the response curves has been changed. Now all standard settings will be regularly taken with either biweekly or monthly cadence. See the calibration plan in the manuals section.
  • A general maintenance of the instrument was carried out on October 13-19, 2021. It involved the maintenance of the camera shutters, the regeneration of the cryostats, the maintenance of the spectrograph slits, and the replacement of the faulty Red exposure meter by a new one.
  • 25/05/2021 User Manual P108 sect. 2.4.1 updated with a table detailing the size of the inter-chip gap in the red arm, since the ETC is not always precise in this regard.
  • Since 2020/09/13 UVES is back online, recommissioned and taking science data as Paranal ramps up to restricted operations status.
  • After 2020/03/24 UVES is shut down as part of the Paranal Safe State due to the COVID-19 pandemic
  • 30/09/2019 For P105 proposals the new P1 online tool and turbulence category constraint are introduced.
  • 01/05/2019 In P103 the calibration programme 4103.L-0942 "Direct Measurement of Atmospheric Dispersion: The long-awaited model validation" will be executed on UVES. All data from calibration programmes become immediately public and can be retrieved via the ESO archive. All calibration programmes since P100 are listed here.
  • 06/10/2017 The second and final part of the intervention on the UVES blue camera lenses, to recover the resolution of the UVES blue arm, is now completed. The resolving power in the UVES blue arm is back to its early-days values, with a most-recent median measurement of R~93,000 (for the 0.4-arcsec slit).
  • 05/02/2017 An intervention to improve the resolution of the blue arm of UVES, which has shown to degrade slowly through the years due to aging of the blue camera, took place. The resolving power measured after the intervention is ~85,000 for the 0.4-arcsec slit. The RMS of the residuals of the wavelength calibration improved by a factor of two.
  • 02/04/2016 An intervention to restore the full performance of the blue arm took place at the period change over to P97. The blue-arm camera is now back to its condition pre-earthquake, i.e., the blue arm delivers a resolving power of ~75,000 for a slit width of 0.4 arcsec.
  • 27/11/2015 A strong earthquake occured on Paranal (http://earthquake.usgs.gov/earthquakes/eventpage/us100041t7#impact_shakemap:us_us100041t7). After optical re-alignment, UVES was still affected. The blue-arm camera suffered from coma which significantly degraded the resolution for slits narrower than 0.9 arcsec. For wider slits, the resolving power was degraded by 5% relative to the offered value. The performance of the red arm kept optimum.
  • 16-31/10/2015 An extra-ordinary maintenance of UVES took place after ten years of (almost) uninterrupted operations.
  • 12-14/05/2015 Malfunctioning and replacement of the Red CCD shutter demanded to re-align and focus the detector unit again.
  • 03-17/04/2015 Improvements of the spectral format in the Blue and Red arms of UVES, including a small rotation of the Red unit, were performed using as reference the 2009 situation so as to minimize differences with the UVES physical model used by the data reduction pipeline. Final re-focussing of the units on April 17 led to optimum resolving power values of R=72,000 and 110,000 on the Blue and Lower/Upper Red CCDs, for slit widths of 0.4 and 0.3 arsec, respectively.
  • 19-23/03/2015 A red-arm pulpo failure occured. The cryostat was dismounted and the shutter replaced. As a result, the spectral format in the UVES red arm changed by 3-5 pixels in the cross-dispersion direction. This was corrected and the unit was re-focussed on March 23.
  • 14/09/2014 On 25 July 2014, a misalignment of the UVES ThAr projection system was found and corrected. This resulted in an artificially high resolution derived from arc frames taken during the first semester of 2014. The resolving power of the instrument was not affected by this problem and therefore remained unchanged.
  • 04/03/2014 UVES echelle (point-like) and slicer reduced data are available through the phase 3 archive query form: http://archive.eso.org/wdb/wdb/adp/phase3_spectral/form. Data is available from year 2000. The current stream of data is processed and published at a roughly monthly cadence.
  • 07/01/2014 UVES red arm CCD shutter failure. After the recovery, the red CCD format resulted shifted by a few px in X,Y with respect to the pre failure situation. This affected mostly cross-disperser #4 settings. This can be seen from the instrument Health Check pages.
  • 01/04/2011 UVES blue cross-disperser (CDs 1 & 2) bearing and worm gear system were upgraded on 2010-11-22. The new mechanism is similar to the one already installed in the red arm in 2005.
  • 04/11/2009 UVES pipeline 4.4.8 supporting new red CCD (and old) data reduction.
  • 01/09/2009 UVES ETC updated with the new QE curve for the Zeus (upgraded CCD).
  • 29/08/2009 UVES documentation updated for P85.
  • 26/08/2009 The UVES red arm quantum efficiency ratio for the new/old MIT CCD is shown below (left panel) as derived in the lab and from standard stars.The EEV (bluer part of the red mosaic) is UNCHANGED.The MIT (redder part of the red mosaic) CCD "Nigel" was replaced with another MIT CCD "Zeus" in early July 2009 with superior Quantum Efficiency above 700-nm. The fringing is also much reduced; Plot (a) shows an extracted FLAMES-UVES flatfield order at about 900-nm with the Old CCD counts multiplied by 2.5 (same integration time). Plot (b) shows one normalised order of a stellar spectrum taken before and after the upgrade using standalone UVES.
  • 15/07/2009: UVES MIT CCD has been replaced and is now back in operation. A report will be made availableshortly. Please note that the current version of the UVES pipeline will not work/give incorrect results for the red arm due to the change in orientation of the CCD. A new version is currently under testing. Please note too that the ETC has notyet been updated with the new efficiencies.
  • 30/06/2009: Minor updates to the UVES user manual for P84 phase II available here (v84.3). Only change compared with v84.2 concerns the TCCD slit viewers.
  • 25/06/2009: Please remember thatthe ETC will only be upgraded with the new sensitivity values for the red-arm after the recommissioning in early July.When the new values are known they will be posted with a link on this webpage. Until then the current ETC should still be used.
  • 26/02/2009: UVES Red Arm will be unavailable for service or visitor mode observations in the period May1st-July 15 due to the upgrade of the red mosaic detector. See Sect.2.2.4 of P84 User Manual for more information.
  • 08/12/2008: UVES ETC for free template: two typos fixed(see here).
  • 22/05/2008: update of the UVES data reduction cookbook for the MIDAS pipeline version 2.9.7(see here).
  • 20/06/2007: the new UVES pipeline version 3.3.1, based on CPL, has been released to support echelle data reduction.
  • 09/02/2007: Ultra Fast Readout mode (Red Arm only, 1x1 binning, low gain) added to the UVES Exposure Time Calculator (v3.1.1).
  • 01/09/2005: A set of eight interference filters for long-slit spectroscopy of extended objects is offered for the first time in P77 (in visitor mode only).
  • 01/06/2005: Starting in P76, the red arm setting with central wavelength 600nm can be used in service mode without the Iodine Cell.
  • 01/06/2005: Due to the unavailability of ared arm setting with central wavelength 760nm, the possibility of usingthe combination of Image Slicer #3 with Dichroic #2 has been implemented.
  • 01/04/2005: Three new standard Dichroic settings with 760 nm have been implemented for use starting in P75: Dic#2 346+760, 390+760 and 437+760.
  • 20/10/2004: The scientific CCD for the UVES blue arm has been exchanged with a new one, called Pavarotti.Information about the new CCD is given in the UVES User Manual (Issue2).
    Both the data obtained with the new and the old CCD can be processed with the new UVES pipeline, version 2.2.0. The latest pipeline version is available here.
  • 01/04/2004: The RRM is offered for UVES in service mode since the beginning of P73. Starting in P75, it is also possible to trigger during visitor mode observations.
    Detailed information on the RRM is given on the USD RRM webpage.
  • 01/04/2004: The format of the Red UVES (and FLAMES-UVES) raw files has changed. The new format is a FITS file with one extension per detector, i.e. the two Red CCD images are now split. The new format cannot be processed with UVES pipeline versions older than 2.1.